OpenStudy (anonymous):

Convert fraction to decimal (or mixed decimal and round to nearest hundredth) 13 6/7

OpenStudy (anonymous):

19.43

OpenStudy (anonymous):

136/7 = (13*7+6)/7 = 97/7 = 13.857 = 13.86

OpenStudy (anonymous):

So which answer is correct

OpenStudy (anonymous):

idk y he did like that